Help Save&Exit Su Gargantic Inc. had net income of $90,000 for 2018; its weighted-average common shares outstanding for 2018 was 18,000 shares. The Companys dividend per share was $0.45, its market price per share was $88, and its book value per share was $76. What was Gargantic incs pric/earnings ratio for 2018? Multiple Choice 176. 5.2 50 None of the other 3 answer choices is the correct price/eamings ratio for Gargantic Inc. for 2018 < Prev 4 of 10辮 Next> 名¢ 0
Answer #1

Price-Earnings Ratio (P/E Ratio)

Price-Earnings Ratio (P/E Ratio) = Market Price of the Share / Earnings per share (EPS)

Earnings per share (EPS) = Net Income / Weighted average common shares outstanding

= $90,000 / 18,000 Shares

= $5.00 per share

Therefore, The Price-Earnings Ratio (P/E Ratio) = Market Price of the Share / Earnings per share (EPS)

= $88 per share / $5.00 per share

= 17.6 Times

“Therefore, The Gargantic Inc’s Price-Earnings Ratio (P/E Ratio) for 2018 = 17.6 Times”